OpenStudy (anonymous):

simplify completely: √24(x^6)

OpenStudy (anonymous):

\[\sqrt{24x^6}\]

OpenStudy (colouringme):

the make it easier, I usually split it --> \[\sqrt{24} \sqrt{x ^{6}}\] try to make factorise it.

OpenStudy (colouringme):

soo.. what's the answer?

OpenStudy (colouringme):

because it's not yet finished, you can simplify it even more! :)

OpenStudy (anonymous):

i have no idea how to do this omg

OpenStudy (colouringme):

okay, factorise 4

OpenStudy (anonymous):

how do you factorise?

OpenStudy (colouringme):

I mean divide, divide with prime numbers

OpenStudy (anonymous):

ohh okay so it's √6 √x^6 ?

OpenStudy (colouringme):

i mean like this, you ignore the roots and x first. So you're left with 24, then do what i did in the attachment. Pay attention as I only use prime numbers to divide it,

OpenStudy (colouringme):

So you got \[2^{3} \times 3\] getting back to the question , put back the roots Now you have this equation \[\sqrt{2^{3} \times 3 \times x ^{6}}\]

OpenStudy (anonymous):

i think i get it now, thank you.

OpenStudy (colouringme):

so the result is?